# Reverse Diffusion Sequential Monte Carlo Samplers

### v1

@@ Appendix D Experiment details
All experiments were conducted on a single NVIDIA A100 GPU with 40 GB GPU memory, and accessed between 16 GB and 32 GB of CPU memory, depending on the task.
D.1 Method and hyperparameter details
We first give an overview of baseline methods we benchmark RDSMC against. We then describe key hyperparameter settings of our method and competing baselines.
D.1.1 Summary of baseline methods
AIS [24].
The annealed importance sampling (AIS) algorithm defines a sequence of geometrically annealed distributions $\rho_{t}(x)$ for $t\in\{0,1,\cdots,T\}$ from an initial proposal $\rho_{0}(x)$ to the desired target $\rho_{T}(x)\propto\pi(x)$ as $\rho_{t}(x)\propto\rho_{0}(x)^{1-\beta_{t}}\pi(x)^{\beta_{t}}$ where $\{\beta_{t}\}_{t=0}^{T}$ is an increasing linear schedule with $\beta_{t}=t/T$. AIS begins by sampling from the initial proposal $\rho_{0}(x)$ and proceeds through a sequence of MCMC transitions, each targeting the intermediate distribution $\rho_{t}(x)$. The resulting weighted samples provide a consistent approximation to the target distribution $\pi(x)$.
SMC [25].
Sequential Monte Carlo (SMC) operates on the same sequence of annealed distributions as AIS but differs by performing resampling at certain iterations. In our implementation, we consider both constant resampling at each step and an adaptive strategy based on an ESS threshold $\kappa_{\mathrm{\text{ess}}}$.
RDMC [8]
Reverse diffusion Monte Carlo (RDMC) builds on the time-reversed Ornstein–Uhlenbeck diffusion:
$dY_{t}=\bigl{\{}Y_{t}+2\,\nabla\log p_{t_{\mathrm{\text{init}}}-t}(Y_{t})\bigr{\}}\,dt+\sqrt{2}\,dB_{t},\quad Y_{0}\sim p_{t_{\mathrm{\text{init}}}}$
@@ D.1.1 Summary of baseline methods
Following the implementation of Grenioux et al. [9], RDMC discretizes the interval $0,t_{\mathrm{\text{init}}}]$ into $T$ steps. At initialization, RDMC employs a Langevin-within-Langevin scheme to sample from $p_{t_{\mathrm{\text{init}}}}$: using Unadjusted Langevin Algorithm (ULA) where the score function is estimated based on DSI. In subsequent steps, RDMC simulates from the reverse SDE where the denoiser $u_{t}(y)$ is obtained by running the Metropolis-adjusted Langevin Algorithm (MALA) targeting $q_{t}(x\mid y)$.
SLIPS [9]
@@
  • The SLIPS initialization also involves a Langevin-within-Langevin procedure. Different from RDMC, SLIPS chooses the initialization time $t_{\mathrm{\text{init}}}$ such that both the marginal $p_{t_{\mathrm{\text{init}}}}(y)$ and the posterior $q_{t_{\mathrm{\text{init}}}}(x\mid y)$ are (approximately) log-concave , which they refer to as the “duality of log-concavity" assumption.
@@
SMS [33].
@@
serves as a non-Markovian denoiser, and one shows that the law of $u_{m}(Y^{1:m})$ converges to $\pi$ in $W_{2}$ at rate $O(\sigma\sqrt{d/m})$ [33].
To sample from $\pi$, one first simulates the entire $M$-tuple $Y^{1:M}$ and then returns $u_{M}(Y^{1:M})$. They employ a ‘Once-At-A-Time’ strategy: draw $Y^{1}$ (whose marginal is $\pi\ \mathcal{N}(0,\sigma^{2}I)$) via Underdamped Langevin, then sequentially sample each $Y^{m}\mid Y^{1:m-1}$ using MCMC targeting the conditional density of $Y^{m}$. Under mild assumptions this sequence of targets becomes increasingly log-concave in $m$, but the initial draw of $Y^{1}$ can be as challenging as sampling from $\pi$ itself when $\sigma$ is large. Although one can estimate scores via importance sampling or inner-loop posterior MCMC, numerical results show a steep degradation in performance unless $\sigma$ is carefully tuned—a manifestation of the same “duality of log-concavity” that forces a trade-off between ease of initialization and overall convergence as reported by Grenioux et al. [48].
D.1.2 Implementation details and hyperparameter settings
Target $n_{\mathrm{is}}$ $n_{\mathrm{\text{steps}}}$ $\delta_{\text{mcmc}}$ $t_{\mathrm{start\text{-}resampling}}$ score est. mcmc kernel
@@ D.1.2 Implementation details and hyperparameter settings
Rings {100} {1, 10} {0.01, 0.05} {0.5} DSI LG
Funnel {100} {1, 10, 50} {0.01, 0.05} {0.5, 0.8} DSI LG
Logistic Reg. { 10, 100} {10, 50} {0.01, 0.005} {0.5, 0.8} TSI LG
Table 3: Hyperparameter grid used for tuning RDSMC. We fix the number of sampling steps at $T=100$, set $m_{\mathrm{\text{steps}}}=1$ MCMC step per AIS transition, and use an ESS threshold of $\kappa_{\mathrm{\text{ess}}}=0.3$. Score estimator and MCMC kernel types are selected via a coarse grid search using validation metrics. RDSMC (IS) and RDSMC (Proposal) use the same grid with $\kappa_{\mathrm{\text{ess}}}=0$, excluding $t_{\mathrm{start\text{-}resampling}}$ as it does apply in these settings.
RDSMC.
We summarize the hyperparameter values used to tune RDSMC and its variants in different experiments in Table˜3. Key hyperparameters are explained below.
We consider the VP diffusion schedule described in Eq.˜16. We fix RDSMC to use $T=100$ time discretization steps evenly spaced between $[0,1]$. The remaining hyperparameters pertain to the MC score estimation and the SMC procedure.
For MC score estimation, we use AIS (Algorithm˜3), with vanilla IS (Algorithm˜2) as a special case when $n_{\mathrm{\text{steps}}}=1$ (in which case $m_{\mathrm{\text{steps}}}$ is not applicable). Asociated hyperparameters include
  • Score estimator type: DSI (Eq.˜5) or TSI (Eq.˜19).
  • MCMC transition kernel: Langevin dynamics (LG) or Hamiltonian Monte Carlo (HMC).
  • $n_{\mathrm{is}}$: Number of importance samples used in Monte Carlo score estimation
  • $n_{\mathrm{\text{steps}}}$: Number of intermediate distributions (i.e., AIS steps).
  • $\delta_{\text{mcmc}}$: Step size used in the MCMC transition kernel.
  • $m_{\mathrm{\text{steps}}}$ : Number of MCMC steps per AIS transition. We use $m_{\mathrm{\text{steps}}}=1$ in all experiments.
For SMC procedures related, we consider
  • $t_{\mathrm{start\text{-}resampling}}$: Starting step for resampling within the SMC trajectory. No resampling is performed for $t>t_{\mathrm{start\text{-}resampling}}$.
  • $\kappa_{\mathrm{\text{ess}}}$: Normalized effective sample size threshold triggering resampling. When $\kappa_{\mathrm{\text{ess}}}=0$, no resampling is applied throughout; this corresponds to RDSMC (IS) when the final-step weighting is used, or RDSMC (Proposal) otherwise. We set $\kappa_{\mathrm{\text{ess}}}=0.3$ in all experiments involving RDSMC.
@@
Table 4: Hyperparameter grid used for tuning baseline methods. AIS, SMC, RDMC, and SLIPS use $T=1,000$ sampling steps, and SMS uses $K=500$ noisy observations. On GMM experiments, we add additional ablation for AIS and SMC with $T=100$ steps.
We select the score estimator and MCMC kernel types through a coarse grid search using validation metrics. And then we primarily tune $n_{\mathrm{is}},n_{\mathrm{\text{steps}}},\delta_{\text{mcmc}}$ and $t_{\mathrm{start\text{-}resampling}}$, as described in Table˜3.
The hyperparameter search grid for baseline methods is provided in Table˜4. We largely follow the implementation of the official codebase of Grenioux et al. [9] (available at https://github.com/h2o64/slips). For the bi-modal GMM experiments, we tune key hyperparameters within a similar range to those used for RDSMC. For the Rings, Funnel, and Bayesian logistic regression tasks, we adopt the settings reported by Grenioux et al. [9].
@@
We restate Assumption A0 from Grenioux et al. [9]
@@
RDMC uses $T=1,000$ discretization steps. To initialize $p_{t_{\mathrm{\text{init}}}}$ in RDMC, the Langevin-within-Langevin algorithm is simulated using 16 MCMC steps and 4 MCMC chains. The chains are initialized with an IS approximation of the posterior powered by 128 particles. The initial sample is drawn from $\mathcal{N}(0,(1-\exp\{-2t_{\mathrm{\text{init}}}\})\mathbb{I}_{d}\})$ and the initial step size is set to $(1-\exp\{-2t_{\mathrm{\text{init}}}\})/2$
@@
SLIPS follows a similar initialization and posterior sampling scheme as RDMC, also with $T=1,000$ discretization steps. However, instead of tuning the value of the initial sampling time $t_{\mathrm{\text{init}}}$, SLIPS makes use of the target variance information and sets $p_{t_{\mathrm{\text{init}}}}(x):=\mathcal{N}(x\mid 0,R^{2}/d+\tau^{2})$.
@@
D.2 Bi-moddal gaussian mixturess
@@
Figure 2: Estimation bias of weight ratio $w_{1}/(w_{1}+w_{2})$ versus dimension $d$. Results are averaged over 5 seeds with error bars showing one standard error. Bias increases with $d$ for all methods, while RDSMC consistently outperforms the baselines.
@@
Figure˜1(b) in the main text displays the comparison among AIS, SMC, RDSMC and its variances for presentation clarity. Here, we present the full results of weight ratio estimation bias including SMS, RDMC, and SLIPS in Figure˜2. We observe that SMS, RDMC, and SLIPS consistently underperform RDSMC across dimensions.
The hyperparameters used for results in Figures˜1(b) and 2 are selected based on minimizing the estimation bias of weight ratio $w_{1}/(w_{1}+w_{2})$. Here, we also report results using hyperparameters tuned by minimizing the $\log Z$ bias in Figure˜3. Results for RDSMC (Proposal), SMS, RDMC, and SLIPS are not included as they do not provide normalization constant estimate. The trends are largely similar, except that the $\log Z$ estimation bias curve appears monotonic.
D.3 Rings and Funnel distributions
@@
The Rings distribution [9] is defined via an inverse polar reparameterization of a base distribution $p_{z}$, which is factorized into two independent univariate marginals $p_{r}$ and $p_{\theta}$:
@@
The Funnel distribution [34] is characterized by a hierarchical density structure in a 10-dimensional space, where the first dimension $x_{1}$ is drawn from a Gaussian distribution with zero mean and variance $\sigma^{2}=9$. The subsequent dimensions $x_{2:10}$ are conditionally Gaussian, with variances exponentially dependent on the first dimension. Formally, the density is
@@
where $\varepsilon>0$ is a regularization hyper-parameter and $H(\pi)=-\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\log\pi(x_{0},x_{1})\mathrm{d}\pi(x_{0},x_{1})$ refers to as the entropy of $\pi$. We evaluate the quality of sampling with regularization $\varepsilon=0.05$ via POT library [49].
@@
where $F_{\mu}$ and $F_{\nu}$ denotes the cumulative distribution function of $\mu$ and $\nu$ respectively. We evaluate the 10-dimensional Funnel samples using the sliced version from [48, Appendix D.1].
@@
For both targets, we use DSI to estimate the score. On Rings, we choose $R=4,\tau=0.15$ that satisfy Assumption˜1 to initialize AIS, SMC, SMS, and SLIPS; and on Funnel, we use $R=2.12,\tau=0.0$, following the implementation of Grenioux et al. [9].
@@
Additional results.
Algorithm Sample W2 Radius TVD Angle TVD $\log Z$ Bias
AIS 0.19 $\pm$ 0.01 0.10 $\pm$ 0.00 0.15 $\pm$ 0.00 0.05 $\pm$ 0.00
SMC 0.19 $\pm$ 0.01 0.10 $\pm$ 0.00 0.15 $\pm$ 0.00 0.05 $\pm$ 0.00
SMS 0.23 $\pm$ 0.01 0.24 $\pm$ 0.01 0.14 $\pm$ 0.00 N/A
RDMC 0.41 $\pm$ 0.01 0.37 $\pm$ 0.00 0.14 $\pm$ 0.00 N/A
SLIPS 0.22 $\pm$ 0.01 0.19 $\pm$ 0.00 0.13 $\pm$ 0.00 N/A
RDSMC 0.26 $\pm$ 0.01 0.13 $\pm$ 0.01 0.21 $\pm$ 0.01 0.03 $\pm$ 0.00
RDSMC(IS) 0.29 $\pm$ 0.01 0.15 $\pm$ 0.01 0.24 $\pm$ 0.00 0.02 $\pm$ 0.01
RDSMC(Proposal) 0.18 $\pm$ 0.01 0.09 $\pm$ 0.00 0.14 $\pm$ 0.00 N/A
Table 5: Additional results for Rings. For all metrics, lower is better.
Additional results for the Rings experiment are provided in Table˜5. We observe that RDSMC (Proposal) achieves the best overall sample quality, as measured by Sample W2, Radius TVD, and Angle TVD. While RDSMC’s Radius TVD closely matches the best performance, its Angle TVD and Sample W2 lag behind other baselines and RDSMC (Proposal). However, both RDSMC and its importance sampling (IS) variant achieve lower bias in estimating the log normalization constant compared to AIS and SMC.
D.4 Bayesian logistic regression

### v2

@@ Appendix E Experiment details
All experiments were conducted on a single NVIDIA A100 GPU with 40 GB GPU memory, and accessed between 16 GB and 32 GB of CPU memory, depending on the task. RDSMC and baseline methods are all implemented in PyTorch [paszke2019pytorch].
E.1 Method and hyperparameter details
We first give an overview of baseline methods we benchmark RDSMC against in §˜E.1.1. We then describe key hyperparameter settings of our method and competing baselines for the main experiments in §˜E.1.2.
E.1.1 Summary of baseline methods
AIS [neal2001annealed].
The annealed importance sampling (AIS) algorithm defines a sequence of geometrically annealed distributions $\rho_{t}(x)$ for $t\in\{0,1,\cdots,T\}$ from an initial proposal $\rho_{0}(x)$ to the desired target $\rho_{T}(x)\propto\pi(x)$ as $\rho_{t}(x)\propto\rho_{0}(x)^{1-\beta_{t}}\pi(x)^{\beta_{t}}$ where $\{\beta_{t}\}_{t=0}^{T}$ is an increasing linear schedule with each $\beta_{t}=t/T$. AIS begins by sampling from the initial proposal $\rho_{0}(x)$ and proceeds through a sequence of MCMC transitions, each targeting the intermediate distribution $\rho_{t}(x)$. The resulting weighted samples provide a consistent approximation to the target distribution $\pi(x)$. The complexity of AIS is $\mathcal{O}(NTn)$ for $N$ final samples, $T$ annealing steps, and $n$ MCMC transitions per step.
SMC [del2006sequential].
Sequential Monte Carlo (SMC) operates on the same sequence of annealed distributions as AIS but differs by performing resampling at certain iterations. In our implementation, we consider both constant resampling at each step and an adaptive strategy based on an ESS threshold $\kappa_{\mathrm{\text{ess}}}$. The complexity of SMC is the same as that of AIS.
RDMC [huang2023reverse].
Reverse diffusion Monte Carlo (RDMC) builds on the time‐reversed Ornstein–Uhlenbeck diffusion:
$dY_{t}=\bigl\{Y_{t}+2\,\nabla\log p_{t_{\mathrm{\text{init}}}-t}(Y_{t})\bigr\}\,dt+\sqrt{2}\,dB_{t},\quad Y_{0}\sim\pi_{t_{\mathrm{\text{init}}}}$
@@ E.1.1 Summary of baseline methods
Following the implementation of grenioux2024stochastic, RDMC discretizes the interval $0,t_{\mathrm{\text{init}}}]$ into $T$ steps. At initialization, RDMC employs a Langevin-within-Langevin scheme to sample from $\pi_{t_{\mathrm{\text{init}}}}$: using ULA where the score function is estimated based on DSI. In subsequent steps, RDMC simulates from the reverse SDE where the denoiser $u_{t}(y)$ is obtained by running MALA targeting $q_{t}(x\mid y)$.
The overall complexity of RDMC is $\mathcal{O}(NM(n_{\textrm{init}}+Tn))$ where $N$ is the number of final samples, $T$ is the number of iterative sampling steps, and the remaining factors are associated with score estimation: $M$ is the number of MCMC chains, while $n_{\textrm{init}}$ and $n$ denote the number of MCMC transition steps in the initial and subsequent stages, respectively.
SLIPS [grenioux2024stochastic].
@@
  • The SLIPS initialization also involves a Langevin-within-Langevin procedure. Different from RDMC, SLIPS chooses the initialization time $t_{\mathrm{\text{init}}}$ such that both the marginal $\pi_{t_{\mathrm{\text{init}}}}(y)$ and the posterior $q_{t_{\mathrm{\text{init}}}}(x\mid y)$ are (approximately) log-concave , which they refer to as the “duality of log-concavity" assumption.
@@
Similar to RDMC, SLIPS’ complexity is $\mathcal{O}(NM(n_{\textrm{init}}+Tn))$ where $N$ is the number of final samples, $T$ is the number of iterative sampling steps, and the remaining factors are associated with score estimation: $M$ is the number of MCMC chains, while $n_{\textrm{init}}$ and $n$ denote the number of MCMC transition steps in the initial and subsequent stages, respectively.
SMS [saremi2023chain].
@@
serves as a non‐Markovian denoiser, and one shows that the law of $u_{m}(Y^{1:m})$ converges to $\pi$ in $W_{2}$ at rate $O(\sigma\sqrt{d/m})$ [saremi2023chain].
To sample from $\pi$, one first simulates the entire $M$-tuple $Y^{1:M}$ and then returns $u_{M}(Y^{1:M})$. They employ a ‘Once-At-A-Time’ strategy: draw $Y^{1}$ via ULA, then sequentially sample each $Y^{m}\mid Y^{1:m-1}$ using MCMC targeting the conditional density of $Y^{m}$. Under mild assumptions this sequence of targets becomes increasingly log-concave in $m$, but the initial draw of $Y^{1}$ can be as challenging as sampling from $\pi$ itself when $\sigma$ is large. Although one can estimate scores via IS or inner-loop posterior MCMC, numerical results show a steep degradation in performance unless $\sigma$ is carefully tuned—a manifestation of the same “duality of log-concavity” that forces a trade-off between ease of initialization and overall convergence as reported by grenioux2023onsampling.
The time complexity of SMS is $O(NTMn)$ for $N$ final samples, $T$ iterative sampling steps, $n$ MCMC transition steps per sampling step, and $M$ importance samples used for score estimation.
E.1.2 Implementation details and hyperparameter settings for the main experiments
Target $M$ $n$ $\delta_{\text{mcmc}}$ $t_{\mathrm{start\text{-}resampling}}$ score est. mcmc kernel
@@ E.1.2 Implementation details and hyperparameter settings for the main experiments
Rings {100} {1, 10} {0.01, 0.05} {0.5} DSI MALA
Funnel {100} {1, 10, 50} {0.01, 0.05} {0.5, 0.8} DSI MALA
Logistic Reg. { 10, 100} {10, 50} {0.01, 0.005} {0.5, 0.8} TSI MALA
Table 3: Hyperparameter grid used for tuning RDSMC in the main experiments. We fix the number of sampling steps at $T=100$, set $l=1$ MCMC step per AIS transition, and use an ESS threshold of $\kappa_{\mathrm{\text{ess}}}=0.3$. Score estimator and MCMC kernel types are selected via a coarse grid search using validation metrics. RDSMC (IS) and RDSMC (Proposal) use the same grid with $\kappa_{\mathrm{\text{ess}}}=0$, excluding $t_{\mathrm{start\text{-}resampling}}$ as it does apply in these settings.
We summarize the hyperparameter values used to tune RDSMC and its variants for the main experiments in Table˜3, where these hyperparameters are explained in §˜B.3.3. The choices of score estimator and MCMC kernel types are determined by a coarse grid search using validation metrics; We primarily tune $M,n,\delta_{\text{mcmc}}$ and $t_{\mathrm{start\text{-}resampling}}$, as described in Table˜3.
@@
Table 4: Hyperparameter grid used for tuning baseline methods in the main experiments. AIS, SMC, RDMC, and SLIPS use $T=1,000$ sampling steps, and SMS uses $K=500$ noisy observations.
The hyperparameter search grid for baseline methods is provided in Table˜4. We largely follow the implementation of the official codebase of grenioux2024stochastic (available at https://github.com/h2o64/slips). For the bi-modal GMM experiments, we tune key hyperparameters within a similar range to those used for RDSMC. For the Rings, Funnel, and Bayesian logistic regression tasks, we adopt the settings reported by grenioux2024stochastic.
@@
We restate Assumption A0 from grenioux2024stochastic.
@@
RDMC uses $T=1,000$ discretization steps. To initialize $\pi_{t_{\mathrm{\text{init}}}}$ in RDMC, the Langevin-within-Langevin algorithm is simulated using 16 MCMC steps and 4 MCMC chains. The chains are initialized with an IS approximation of the posterior powered by 128 particles. The initial sample is drawn from $\mathcal{N}(0,(1-\exp\{-2t_{\mathrm{\text{init}}}\})\mathbb{I}_{d}\})$ and the initial step size is set to $(1-\exp\{-2t_{\mathrm{\text{init}}}\})/2$
@@
SLIPS follows a similar initialization and posterior sampling scheme as RDMC, also with $T=1,000$ discretization steps. However, instead of tuning the value of the initial sampling time $t_{\mathrm{\text{init}}}$, SLIPS makes use of the target variance information and sets $\pi_{t_{\mathrm{\text{init}}}}(x):=\mathcal{N}(x\mid 0,R^{2}/d+\tau^{2})$.
@@
E.2 Bi-modal gaussian mixtures
@@
The hyperparameters used for results in Figure˜1(b) are selected based on minimizing the estimation bias of weight ratio $w_{1}/(w_{1}+w_{2})$. Here, we also report results using hyperparameters tuned by minimizing the $\log Z$ bias in Figure˜2. Results for RDSMC (Proposal), SMS, RDMC, and SLIPS are not included as they do not provide normalization constant estimate. The trends are largely similar, except that the $\log Z$ estimation bias curve appears monotonic.
E.3 Rings and Funnel distributions
@@
The Rings distribution [grenioux2024stochastic] is defined via an inverse polar reparameterization of a base distribution $p_{z}$, which is factorized into two independent univariate marginals $p_{r}$ and $p_{\theta}$:
@@
The Funnel distribution [neal2003slice] is characterized by a hierarchical density structure in a 10-dimensional space, where the first dimension $x_{1}$ is drawn from a Gaussian distribution with zero mean and variance $\sigma^{2}=9$. The subsequent dimensions $x_{2:10}$ are conditionally Gaussian, with variances exponentially dependent on the first dimension. We follow grenioux2024stochastic to use the following Funnel density
@@
where $\varepsilon>0$ is a regularization hyper-parameter and $H(\pi)=-\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\log\pi(x_{0},x_{1})\mathrm{d}\pi(x_{0},x_{1})$ refers to as the entropy of $\pi$. We evaluate the quality of sampling with regularization $\varepsilon=0.05$ via POT library [flamary2021pot].
@@
where $F_{\mu}$ and $F_{\nu}$ denotes the cumulative distribution function of $\mu$ and $\nu$ respectively. We evaluate the 10-dimensional Funnel samples using the sliced version from [grenioux2023onsampling, Appendix D.1].
@@
For both targets, we use DSI to estimate the score. On Rings, we choose $R=4,\tau=0.15$ that satisfy Assumption˜1 to initialize AIS, SMC, SMS, and SLIPS; and on Funnel, we use $R=2.12,\tau=0.0$, following the implementation of grenioux2024stochastic.
@@
E.4 Bayesian logistic regression